2. The following table lists the nominal federal spending on nondefense-related education and

training (in billion dollars) and CPI (price index) in 1962, 1965, 2001, and 2004.

2. The following table lists the nominal federal spending on nondefense-related education and training (in billion dollars) and CPI (price index) in 1962, 1965, 2001, and 2004. 1962 1965 2001 2004 Nominal Spending $9.6 $19.5 $178.4 $217.5 30 31.2 175.1 185.2 CPI (a) Compare the change in nominal spending from 1962 to 1965 and that from 2001 to 2004, which increase was larger? (b) Compare the change in real spending from 1962 to 1965 and that from 2001 to 2004, which increase was larger? Note that you need to calculate the real spending first Show the detail of your calculation.
Answer #1

a) Change in nominal spending from 1962 to 1965 = (Spending in 1965-Spending in 1962)/Spending in 1962
= (19.5-9.6)/9.6 = 9.9/9.6 = 1.03125
So, Change in nominal spending from 1962 to 1965 is 103.125%. (Converting in percentage form)

Change in nominal spending from 2001 to 2004 = (Spending in 2004-Spending in 2001)/Spending in 2001
= (217.5-178.4)/178.4 = 39.1/178.4 = 0.2192
So, Change in nominal spending from 2001 to 2004 is 21.92%. (Converting in percentage form)

Change in nominal spending from 1962 to 1965 is larger.

b) Firstly we find real spending
Real spending in 1962 = nominal spending from 1962/CPI in 1962 = 9.6/30 = 0.32
Real spending in 1965 = nominal spending from 1965/CPI in 1965 = 19.5/31.2 = 0.625
Real spending in 2001 = nominal spending from 2001/CPI in 2001 = 178.4/175.1 = 1.02
Real spending in 2004 = nominal spending from 2004/CPI in 2004 = 217.5/185.2 = 1.17

Change in real spending from 1962 to 1965 = (Spending in 1965-Spending in 1962)/Spending in 1962
= (0.625-0.32)/0.32 = 0.305/0.32 = 0.9531
So, Change in real spending from 1962 to 1965 is 95.31%. (Converting in percentage form)

Change in real spending from 2001 to 2004 = (Spending in 2004-Spending in 2001)/Spending in 2001
= (1.17-1.02)/1.02 = 0.15/1.02 = 0.1471
So, Change in real spending from 2001 to 2004 is 14.71%. (Converting in percentage form)

Change in real spending from 1962 to 1965 is larger.